The Bold One is 3 times The Timid One's age. in 12 years, The Bold One will be 2 times the Timid One's age. How old is the Timid One now?
Answer by scott8148(6628)

"The Bold One is 3 times The Timid One's age" __ B=3T

"in 12 years, The Bold One will be 2 times the Timid One's age" __ B+12=2(T+12) __ B+12=2T+24

substituting __ 3T+12=2T+24

subtracting 2T+12 __ T=12